OpenStudy (anonymous):

the gravitational force, F, on a rocket at a distance, from the center of the earth is given by F = k / (r^2), where k = 10 ^3 newton*km^2. when the rocket is 10^4 km from the center of the earth, its moving away at .2 km/sec. how fast is the gravitational force changing at that moment? therefore what is the rate of change of the gravitational force in newton/ second?

OpenStudy (amistre64):

rate of change tells me this thing wants derivatives of associated equations

OpenStudy (amistre64):

the gravitational force, F, on a rocket at a distance, r from the center of the earth is given by: F = k / (r^2), where k = 10 ^3 newton*km^2. when the rocket is 10^4 km from the center of the earth, its moving away at .2 km/sec. how fast is the gravitational force changing at that moment? therefore what is the rate of change of the gravitational force in newton/ second? hmm, might be good to draw a picture to make sure we got all the parts

OpenStudy (amistre64):



OpenStudy (amistre64):

all I can tell is that we should take the derivative of F F'(r) = (-2)10^3 r^-3 r'

OpenStudy (amistre64):

r' at that moment is .2 soo i would take a gander and say: F'(10^3) = -2(10^4)^-3 (10^3) (.2) \[F'(10^3)=-2(.2)\frac{10^3}{10^{12}}=-2(.2)(10)^{-9}\]

OpenStudy (amistre64):

F = -4.0 × 10^-10 with any luck; but I cant be too certain that I even got the question right to begin with

OpenStudy (anonymous):

you were right with what the question was asking!

OpenStudy (amistre64):

yay!! one of my personalities must be a genuis :)

OpenStudy (anonymous):

haha yes you are my savior! but wouldn't you plug in .2 as f'x

OpenStudy (amistre64):

.2 is the rate of change of the radius; r'

OpenStudy (anonymous):

because the answer is apparently wrong :(

OpenStudy (amistre64):

by deriving f(r) we get f'(r)* r'

OpenStudy (anonymous):

i get that F'(r) = (-2)10^3 r^-3 r', but where did the -2(10^4)^-3 (10^3) (.2) come from?

OpenStudy (amistre64):

i might have transposed some things; let me see

OpenStudy (amistre64):

\[F(r)=kr(t)^{-2}\] \[F'(r)=-2kr(t)^{-3}*r'(t)\] \[F'(10^4)=-2k(10^4)^{-3}*.2\] \[F'(10^4)=-2(10^3)(10^4)^{-3}*2(10^{-1})\] \[F'(10^4)=-4x(10^3)(10^4)^{-3}(10^{-1})\] \[F'(10^4)=-4x(10^2)(10^{-12})\] \[F'(10^4)=-4\ x\ 10^{-10}\] does that ring more true?

OpenStudy (anonymous):

OMG i am sooooo sorry i it's 10 ^ 13 on the question not 10 ^3... you helped me so much but i told you the wrong thing :((((((((

OpenStudy (amistre64):

:) then with any luck the concept is sound, just recalculate with the correct value and see if we are good

OpenStudy (anonymous):

im so sorry and thank you for helping me i was so stuck! and yes i followed your steps and it was the correct answer!! could you be so kind to help me with my other questions? i really have no luck with optimization problems

OpenStudy (amistre64):

is that the spreading oil tanker one?

OpenStudy (anonymous):

yes the gasoline one: Gasoline is pouring into a cylindrical tank of radius 4 feet. When the depth of the gasoline is 6 feet, the depth is increasing at .4 ft/sec. How fast is the volume of gasoline changing at that instant? Round your answer to three decimal places.